# Interpretable Concept-Based Memory Reasoning

### v1

@@ 1 Introduction
The opaque decision process of deep learning (DL) systems represents one of the most fundamental problems in modern artificial intelligence (AI). For this reason, eXplainable AI (XAI) [1, 2, 3] is currently one of the most active research areas in AI. Among XAI techniques, Concept Bottleneck Models (CBMs) [4, 5, 6, 7, 8] represented a significant innovation that made DL models explainable by design by introducing a layer of human-interpretable concepts within DL architectures. In CBMs, first, a concept encoder maps low-level raw features (e.g., an image’s pixels) to high-level interpretable concepts (e.g., “red” and “round”), then a task predictor uses the learned concepts to solve a downstream task (e.g., “apple”). This way, each task prediction can be traced back to a specific pattern of concepts, thus allowing CBMs to provide explanations in terms of high-level interpretable concepts (e.g., concepts “red” and “round” were both active when the model classified an image as “apple”) rather than low-level raw features (e.g., there were 100 red pixels when the model classified an image as “apple”). In other terms, a CBM’s task predictor allows understanding what the model is seeing in a given input rather than simply pointing to where it is looking [9].
@@ 2 Preliminary
[4, 10] are functions composed of (i) a concept encoder $g:X\to C$ mapping each entity $x\in X\subseteq\mathbb{R}^{d}$ (e.g., an image) to a set of $n_{C}$ concepts $c\in C$ (e.g., “red”, “round”), and (ii) a task predictor $f:C\to Y$ mapping concepts to the class $y\in Y$ (e.g., “apple”) representing a downstream task. For simplicity, in this paper, a single task class is discussed, as multiple tasks can be encoded by instantiating multiple task predictors. When sigmoid activations are used for concepts and task predictions, we can consider $g$ and $f$ as parameterizing a Bernoulli distribution of truth assignments to propositional boolean concepts and tasks. For example, $g_{red}(x)=0.8$ means that there is an 80% probability for the proposition “$x$ is red" to be true. During training, concept and class predictions $(c,y)$ are aligned with ground-truth labels $(\hat{c},\hat{y})$. This architecture and training allows CBMs to provide explanations for class predictions indicating the presence or absence of concepts. Another main advantage of these models is that, at test time, human experts may also intervene on mispredicted concept labels to improve CBMs’ task performance and extract counterfactual explanations [4, 11]. However, the task prediction $f$ is still often a black-box model to guarantee high performances, thus not providing any insight into which concepts are used and how they are composed to reach the final prediction.
@@ 3.1.1 CMR rule selector
We model the rule selector $p(r|x)$ as a mixture of ${n_{R}}\in\mathbb{N}$ rule distributions. The selector "selects" a rule from a set of rule distributions (i.e. the components of the mixture), which we call a rulebook. The rulebook is jointly learned and can be inspected and verified at every stage of the learning. Architecturally, this is akin to an attention mechanism over a differentiable memory [12].
@@ 4.1 Expressivity
CMR is a universal binary classifier [13] if $n_{R}\geq 3$.
@@ 4.2 Interpretability
Prototype-based rules. In order to develop interpretable rules, we focus on standard theories in cognitive science [14] by looking at those rules that are prototypical of the concept activations on which they are applied to. Prototype-based models are often considered one of the main categories of interpretable models [9] and have been investigated in the context of concept-based models [15]. However, CMR distinguishes itself from prototype-based models in two significant ways. First, in prototypical concept-based models, prototypes are built directly in the input space, such as images [16] or on part of it [17], and are often used to automatically build concepts. However, this approach carries similar issues w.r.t. traditional input-based explanations, like saliency maps (i.e. a prototype made by a strange pattern of pixels can still remain unclear to the user). In contrast, in CMR, prototypes (i.e. rules) are built on top of human-understandable concepts and are therefore interpretable as well. Second, differently from prototype-based models, CMR assigns a logical interpretation to prototypes as conjunctive rules. Unlike prototype networks that assign class labels only based on proximity to a prototype, CMR determines class labels through the symbolic evaluation of prototype rules on concepts. Therefore, prototypes should be representative of concept activations but also provide a correct task prediction. This double nature of CMR prototypes, as both representative of concept activations and conjunctive rules for task prediction, imposes an additional constraint. In particular, only prototypes of positive task instances (i.e. $y=1$) can behave as good classification rules. The conjunctive nature of the rule makes that a positive evaluation is only possible if all (relevant) conjuncts match with the corresponding concept predictions. In contrast, a negative evaluation can simply be obtained if at least one of the conjuncts does not match with the concept predictions. Therefore, the rules selected for negative instances ($y=0$) will not be required to be representative of the corresponding concept predictions.
@@
Interventions. In contrast to existing CBMs, which only allow prediction-time concept interventions, the global interpretability of CMR allows humans to intervene also on the rules that will be exploited for task prediction. These interactions may occur during the training phase to directly shape the model that is being learned, and may come in various forms. A first approach involves the manual inclusion of rules into the rule book, thereby integrating expert knowledge into the model [18]. The selector mechanism within the model learns to choose among the rules acquired through training and those manually incorporated. As a result, the rules acquired through training can change after adding expert rules. A second approach consists of modifying the learned rules themselves, by altering the role $r_{i}$ of a concept within a rule. For instance, setting the logit of $P_{i}$ to 0 ensures that $c_{i}$ cannot exhibit a positive role in that rule, and setting the logits of both $P_{i}$ and $N_{i}$ to 0 ensures irrelevance. This type of intervention could be exploited to remove (or prevent) biases and ensure counterfactual fairness [19].
@@ 4.3 Verification
One of the main properties of CMR is that, at decision time, it explicitly represents the task prediction as a set of conjunctive rules. Logically, the mixture semantics of the selector can be interpreted as a disjunction, leading to a standard semantics in stochastic logic programming [20, 21]. As the only neural component is in the selection, task predictions generated using CMR’s global formula (cf. Section 4.2) can be automatically verified by using standard tools of formal verification (e.g. model checking), no matter which rule will be selected. Being able to verify properties prior to deployment of the model strongly sets CMR apart from existing models, where verification tasks can only be applied at prediction time. In particular, given any propositional logical formula $\alpha$ over the propositional language $\{c_{1},c_{2},...,c_{n_{C}}\}\cup\{r^{\smash{(j)}}_{i}\}$, $\alpha$ can be automatically verified to logically follow from Equation 1. In logical terms, if the formula $\alpha$ is entailed by Equation 1, it means that $\alpha$ must be true every time Equation 1 is used for prediction.
@@ 5 Learning problem
The maximum likelihood approach only focuses on the prediction accuracy of the model. However, as discussed in Section 4.2, we look for the set of learned rules $r$ to represent good prototypes of concept predictions, as in prototype-based learning [15]. To drive the learning of representative positive prototypes when we observe a positive value for the task, i.e. when $y=1$, we add a regularization term to the objective. Intuitively, every time a rule is selected for a given input instance $x$ with task label $y=1$, we want the rule to be as close as possible to the observed concept prediction. At the same time, since the number of rules is limited and the possible concept activations are combinatorial, the same rule is expected to be selected for different concept activations. When this happens, we will favor rules that assign an irrelevant role to the inconsistent concepts in the activations. The regularized objective is:
@@ 6.1 Experimental setting
Data & task setup. We base our experiments on four different datasets commonly used to evaluate CBMs: MNIST+ [22], where the task is to predict the sum of two digits; C-MNIST, where we adapted MNIST to the task of predicting whether a colored digit is even or odd; MNIST+, where we removed the concepts for the digits 0 and 1 from the concept set; CelebA [23], a large-scale face attributes dataset with more than 200K celebrity images, each with 40 concept annotations [^3] We remove the concepts $Wavy\_Hair$, $Black\_Hair$ and $Male$ from the concept set and instead use them as tasks.; CUB [24], where the task is to predict bird species based on bird characteristics; and CEBAB [25], a text-based task where reviews are classified as positive or negative based on different criteria (e.g. food, ambience, service, etc). These datasets range across different concept set quality, i.e. complete (MNIST+, C-MNIST, CUB) vs incomplete (CelebA, MNIST+), and different complexities of the concept prediction task, i.e. easy (MNIST+, MNIST+, C-MNIST), medium (CEBAB) and hard (CelebA, CUB).
Evaluation. To measure classification performance on tasks and concepts, we compute subset accuracy and regular accuracy, respectively. For CUB, we instead compute the Area Under the Receiver Operating Characteristic Curve [26] for the tasks due to the large class imbalance. All metrics are reported using the mean and the standard error of the mean over three different runs with different initializations.
Baselines. In our experiments, we compare CMR with existing CBM architectures. We consider Concept Bottleneck Models with different task predictors: linear, multi-layer (MLP), decision-tree (DT) and XGBoost (XG). Moreover, we add two state-of-the-art CBMs: Concept Embedding Models (CEM) [27] and Deep Concept Reasoner (DCR) [11]. We employ hard concepts in CMR and our competitors, avoiding the problem of input distribution leakage that can affect task accuracy [28, 29] (see Appendix C for additional details).
@@ 7 Related works
In recent years, XAI techniques have been criticized for their vulnerability to data modifications [30, 31], insensitivity to reparameterizations, [32], and lacking meaningful interpretations for non-expert users [33]. To address these issues, Concept-based methods [34, 35, 5, 10] have emerged, offering explanations in terms of human-interpretable features, a.k.a. concepts. Concept Bottleneck Models [4] take a step further by directly integrating these concepts as explicit intermediate network representations. Concept Embeddings Models (CEMs) [7, 8, 11] close the accuracy gap with black-box models through vectorial concept representations. However, they still harm the final interpretability, as it is unclear what information is contained in the embeddings. In contrast, CMR closes the accuracy gap by exploiting a neural rule selector coupled with learned symbolic logic rules. As a result, CMR’s task prediction is fully transparent, allowing experts to see how concepts are being used for task prediction, and allowing intervention and automatic verification of desired properties. To the best of our knowledge, there is only one other attempt at analyzing CBMs’ task prediction in terms of logical formulae, namely DCR [11]. For a given example, DCR predicts and subsequently evaluates a (fuzzy) logic rule. As rules are predicted on a per-example basis, the global behavior of DCR cannot be inspected, rendering interaction (e.g. manually adding expert rules) and verification impossible. In contrast, CMR learns (probabilistic) logic rules in a memory, allowing for inspection, interaction and verification.
The use of logic rules by CMR for interpretability purposes aligns it closely with the field of neurosymbolic AI [36, 37]. Here, logic rules [38, 39, 18] or logic programs [40, 22, 21] are used in combination with neural networks through the use of neural predicates [22]. Concepts in CMR are akin to a propositional version of neural predicates. However, in CMR, the set of rules is learned instead of given by the human and concept direct supervision is used for human alignment.
@@ C.1 Datasets
This dataset [22] consists of pairs of images, where each image is an MNIST image of a digit and the task is the sum of the two digits. For these tasks, all concepts are relevant. There is a total of 30,000 training examples and 5,000 test examples.
@@
We derive this dataset from MNIST [41], taking the MNIST training (60,000 examples) and test set (10,000 examples), randomly coloring each digit either red or green, and adding these two colors as concepts. There are two tasks: The first is whether the digit is even, and the second is whether it is odd. For these tasks, the concepts related to color are irrelevant.
@@
This is a large-scale face attributes dataset with more than 200K celebrity images [23]. Each image has 40 concept annotations. As tasks, we take the concepts $Wavy\_Hair$, $Black\_Hair$, and $Male$, removing them from the concept set.
@@
In this dataset, the task is to predict bird species from images, where the concepts are 112 bird characteristics such as tail color, wing pattern, etc  [24]. CUB originally consists of 200 tasks, of which we take the first 10. Some concepts are strongly imbalanced (some are True in only ±$\pm 0.5\%$ of examples, others in $40\%$, etc.) and there is a large task imbalance (each task is True in only ±$\pm 0.5\%$ of examples).
@@
This is a text-based dataset that consists of reviews, where the task is to classify them as positive or negative [25]. There are 12 concepts (food, ambience, service, noise, each either unknown, bad or good) and 2 tasks (positive or negative review).
@@ C.2 Training
When departing from pure probabilistic semantics, CBMs allow not only the use of concepts as binary variables, but they allow for concepts to be passed to the task predictor together with their prediction scores, which is called employing soft concepts (vs hard concepts). While some CBMs use soft concepts as this results in higher task accuracy, the downside of this is that the use of soft concepts also comes with the introduction of input distribution leakage: The concept probabilities encode much more information than what is related to the concepts, severely harming the interpretability of the model [28, 29]. For this reason, in our experiments, all models use hard concepts, which is realized by thresholding the soft concept predictions at 50%.
@@
For MNIST+, MNIST+ and C-MNIST, we train directly on the images. For CelebA and CUB, instead of training on the images, we train the models on pretrained ResNet18 embeddings [42]. Specifically, using the torchvision library, we first resize the images to width and height 224 (using bi-linear interpolation), then normalize them per channel with means $(0.485,0.456,0.406)$ and standard-deviations $(0.229,0.224,0.225)$ (for CelebA only). Finally, we remove the last (classification) layer from the pretrained ResNet18 model, use the resulting model on each image, and flatten the output, resulting in an embedding. For CEBAB, we use a pretrained BERT model [43] to transform the input into embeddings. Specifically, using the transformers library [44], we create a BERT model for sequence classification from the pretrained model ’bert-base-uncased’ with 13 labels (1 per concept and 1 representing both tasks). Then, we finetune this model for 10 epochs with batch size 2, 500 warmup steps, weight decay 0.01 and 8 gradient accumulation steps. After training, we use this model to transform each example into an embedding by outputting the last hidden states for that example.

### v2

@@ 1 Introduction
The opaque decision process of deep learning (DL) systems represents one of the most fundamental problems in modern artificial intelligence (AI). For this reason, eXplainable AI (XAI) [guidotti2018survey, adadi2018peeking, arrieta2020explainable] is currently one of the most active research areas in AI. Among XAI techniques, Concept Bottleneck Models (CBMs) [koh2020concept, alvarez2018towards, chen2020concept, zarlenga2023towards, kim2023probabilistic] represented a significant innovation that made DL models explainable by design by introducing a layer of human-interpretable concepts within DL architectures. In CBMs, first, a concept encoder maps low-level raw features (e.g., an image’s pixels) to high-level interpretable concepts (e.g., “red” and “round”), then a task predictor uses the learned concepts to solve a downstream task (e.g., “apple”). This way, each task prediction can be traced back to a specific pattern of concepts, thus allowing CBMs to provide explanations in terms of high-level interpretable concepts (e.g., concepts “red” and “round” were both active when the model classified an image as “apple”) rather than low-level raw features (e.g., there were 100 red pixels when the model classified an image as “apple”). In other terms, a CBM’s task predictor allows understanding what the model is seeing in a given input rather than simply pointing to where it is looking [rudin2019stop].
@@ 2 Preliminary
[koh2020concept, poeta2023concept] are functions composed of (i) a concept encoder $g:X\to C$ mapping each entity $x\in X\subseteq\mathbb{R}^{d}$ (e.g., an image) to a set of $n_{C}$ concepts $c\in C$ (e.g., “red”, “round”), and (ii) a task predictor $f:C\to Y$ mapping concepts to the class $y\in Y$ (e.g., “apple”) representing a downstream task. For simplicity, in this paper, a single task class is discussed, as multiple tasks can be encoded by instantiating multiple task predictors. When sigmoid activations are used for concepts and task predictions, we can consider $g$ and $f$ as parameterizing a Bernoulli distribution of truth assignments to propositional boolean concepts and tasks. For example, $g_{red}(x)=0.8$ means that there is an 80% probability for the proposition “$x$ is red" to be true. During training, concept and class predictions $(c,y)$ are aligned with ground-truth labels $(\hat{c},\hat{y})$. This architecture and training allows CBMs to provide explanations for class predictions indicating the presence or absence of concepts. Another main advantage of these models is that, at test time, human experts may also intervene on mispredicted concept labels to improve CBMs’ task performance and extract counterfactual explanations [koh2020concept, barbiero2023interpretable]. However, the task prediction $f$ is still often a black-box model to guarantee high performances, thus not providing any insight into which concepts are used and how they are composed to reach the final prediction.
@@ 3.1.1 CMR rule selector
We model the rule selector $p(r|x)$ as a mixture of ${n_{R}}\in\mathbb{N}$ rule distributions. The selector "selects" a rule from a set of rule distributions (i.e. the components of the mixture), which we call a rulebook. The rulebook is jointly learned and can be inspected and verified at every stage of the learning. Architecturally, this is akin to an attention mechanism over a differentiable memory [weston2014memory].
@@ 4.1 Expressivity
CMR is a universal binary classifier [hornik1989multilayer] if $n_{R}\geq 3$.
@@ 4.2 Interpretability
Prototype-based rules. In order to develop interpretable rules, we focus on standard theories in cognitive science [rosch1978principles] by looking at those rules that are prototypical of the concept activations on which they are applied to. Prototype-based models are often considered one of the main categories of interpretable models [rudin2019stop] and have been investigated in the context of concept-based models [rudin2022interpretable]. However, CMR distinguishes itself from prototype-based models in two significant ways. First, in prototypical concept-based models, prototypes are built directly in the input space, such as images [li2018deep] or on part of it [chen2019looks], and are often used to automatically build concepts. However, this approach carries similar issues w.r.t. traditional input-based explanations, like saliency maps (i.e. a prototype made by a strange pattern of pixels can still remain unclear to the user). In contrast, in CMR, prototypes (i.e. rules) are built on top of human-understandable concepts and are therefore interpretable as well. Second, differently from prototype-based models, CMR assigns a logical interpretation to prototypes as conjunctive rules. Unlike prototype networks that assign class labels only based on proximity to a prototype, CMR determines class labels through the symbolic evaluation of prototype rules on concepts. Therefore, prototypes should be representative of concept activations but also provide a correct task prediction. This double nature of CMR prototypes, as both representative of concept activations and conjunctive rules for task prediction, imposes an additional constraint. In particular, only prototypes of positive task instances (i.e. $y=1$) can behave as good classification rules. The conjunctive nature of the rule makes that a positive evaluation is only possible if all (relevant) conjuncts match with the corresponding concept predictions. In contrast, a negative evaluation can simply be obtained if at least one of the conjuncts does not match with the concept predictions. Therefore, the rules selected for negative instances ($y=0$) will not be required to be representative of the corresponding concept predictions.
@@
Interventions. In contrast to existing CBMs, which only allow prediction-time concept interventions, the global interpretability of CMR allows humans to intervene also on the rules that will be exploited for task prediction. These interactions may occur during the training phase to directly shape the model that is being learned, and may come in various forms. A first approach involves the manual inclusion of rules into the rule book, thereby integrating expert knowledge into the model [diligenti2017semantic]. The selector mechanism within the model learns to choose among the rules acquired through training and those manually incorporated. As a result, the rules acquired through training can change after adding expert rules. A second approach consists of modifying the learned rules themselves, by altering the role $r_{i}$ of a concept within a rule. For instance, setting the logit of $P_{i}$ to 0 ensures that $c_{i}$ cannot exhibit a positive role in that rule, and setting the logits of both $P_{i}$ and $N_{i}$ to 0 ensures irrelevance. This type of intervention could be exploited to remove (or prevent) biases and ensure counterfactual fairness [Kusner2017].
@@ 4.3 Verification
One of the main properties of CMR is that, at decision time, it explicitly represents the task prediction as a set of conjunctive rules. Logically, the mixture semantics of the selector can be interpreted as a disjunction, leading to a standard semantics in stochastic logic programming [cussens2001parameter, Winters_Marra_Manhaeve_Raedt_2022]. As the only neural component is in the selection, task predictions generated using CMR’s global formula (cf. Section 4.2) can be automatically verified by using standard tools of formal verification (e.g. model checking), no matter which rule will be selected. Being able to verify properties prior to deployment of the model strongly sets CMR apart from existing models, where verification tasks can only be applied at prediction time. In particular, given any propositional logical formula $\alpha$ over the propositional language $\{c_{1},c_{2},...,c_{n_{C}}\}\cup\{r^{\smash{(j)}}_{i}\}$, $\alpha$ can be automatically verified to logically follow from Equation 1. In logical terms, if the formula $\alpha$ is entailed by Equation 1, it means that $\alpha$ must be true every time Equation 1 is used for prediction.
@@ 5 Learning problem
The maximum likelihood approach only focuses on the prediction accuracy of the model. However, as discussed in Section 4.2, we look for the set of learned rules $r$ to represent good prototypes of concept predictions, as in prototype-based learning [rudin2022interpretable]. To drive the learning of representative positive prototypes when we observe a positive value for the task, i.e. when $y=1$, we add a regularization term to the objective. Intuitively, every time a rule is selected for a given input instance $x$ with task label $y=1$, we want the rule to be as close as possible to the observed concept prediction. At the same time, since the number of rules is limited and the possible concept activations are combinatorial, the same rule is expected to be selected for different concept activations. When this happens, we will favor rules that assign an irrelevant role to the inconsistent concepts in the activations. The regularized objective is:
@@ 6.1 Experimental setting
Data & task setup. We base our experiments on four different datasets commonly used to evaluate CBMs: MNIST+ [manhaeve2018deepproblog], where the task is to predict the sum of two digits; C-MNIST, where we adapted MNIST to the task of predicting whether a colored digit is even or odd; MNIST+, where we removed the concepts for the digits 0 and 1 from the concept set; CelebA [liu2015faceattributes], a large-scale face attributes dataset with more than 200K celebrity images, each with 40 concept annotations [^3] We remove the concepts $Wavy\_Hair$, $Black\_Hair$ and $Male$ from the concept set and instead use them as tasks.; CUB [cub], where the task is to predict bird species based on bird characteristics; and CEBAB [abraham2022cebab], a text-based task where reviews are classified as positive or negative based on different criteria (e.g. food, ambience, service, etc). These datasets range across different concept set quality, i.e. complete (MNIST+, C-MNIST, CUB) vs incomplete (CelebA, MNIST+), and different complexities of the concept prediction task, i.e. easy (MNIST+, MNIST+, C-MNIST), medium (CEBAB) and hard (CelebA, CUB).
Evaluation. To measure classification performance on tasks and concepts, we compute subset accuracy and regular accuracy, respectively. For CUB, we instead compute the Area Under the Receiver Operating Characteristic Curve [10.1023/A:1010920819831] for the tasks due to the large class imbalance. All metrics are reported using the mean and the standard error of the mean over three different runs with different initializations.
Baselines. In our experiments, we compare CMR with existing CBM architectures. We consider Concept Bottleneck Models with different task predictors: linear, multi-layer (MLP), decision-tree (DT) and XGBoost (XG). Moreover, we add two state-of-the-art CBMs: Concept Embedding Models (CEM) [EspinosaZarlenga2022cem] and Deep Concept Reasoner (DCR) [barbiero2023interpretable]. We employ hard concepts in CMR and our competitors, avoiding the problem of input distribution leakage that can affect task accuracy [marconato2022glancenets, havasi2022addressing] (see Appendix C for additional details).
@@ 7 Related works
In recent years, XAI techniques have been criticized for their vulnerability to data modifications [kindermans2019reliability, ghorbani2019interpretation], insensitivity to reparameterizations, [adebayo2018sanity], and lacking meaningful interpretations for non-expert users [poursabzi2021manipulating]. To address these issues, Concept-based methods [kim2018interpretability, ghorbani2019towards, alvarez2018towards, poeta2023concept] have emerged, offering explanations in terms of human-interpretable features, a.k.a. concepts. Concept Bottleneck Models [koh2020concept] take a step further by directly integrating these concepts as explicit intermediate network representations. Concept Embeddings Models (CEMs) [zarlenga2023towards, kim2023probabilistic, barbiero2023interpretable] close the accuracy gap with black-box models through vectorial concept representations. However, they still harm the final interpretability, as it is unclear what information is contained in the embeddings. In contrast, CMR closes the accuracy gap by exploiting a neural rule selector coupled with learned symbolic logic rules. As a result, CMR’s task prediction is fully transparent, allowing experts to see how concepts are being used for task prediction, and allowing intervention and automatic verification of desired properties. To the best of our knowledge, there is only one other attempt at analyzing CBMs’ task prediction in terms of logical formulae, namely DCR [barbiero2023interpretable]. For a given example, DCR predicts and subsequently evaluates a (fuzzy) logic rule. As rules are predicted on a per-example basis, the global behavior of DCR cannot be inspected, rendering interaction (e.g. manually adding expert rules) and verification impossible. In contrast, CMR learns (probabilistic) logic rules in a memory, allowing for inspection, interaction and verification.
The use of logic rules by CMR for interpretability purposes aligns it closely with the field of neurosymbolic AI [garcez2022neural, marra2024statistical]. Here, logic rules [xu2018semantic, badreddine2022logic, diligenti2017semantic] or logic programs [sourek2018lifted, manhaeve2018deepproblog, Winters_Marra_Manhaeve_Raedt_2022] are used in combination with neural networks through the use of neural predicates [manhaeve2018deepproblog]. Concepts in CMR are akin to a propositional version of neural predicates. However, in CMR, the set of rules is learned instead of given by the human and concept direct supervision is used for human alignment.
@@ C.1 Datasets
This dataset [manhaeve2018deepproblog] consists of pairs of images, where each image is an MNIST image of a digit and the task is the sum of the two digits. For these tasks, all concepts are relevant. There is a total of 30,000 training examples and 5,000 test examples.
@@
We derive this dataset from MNIST [mnistLecun], taking the MNIST training (60,000 examples) and test set (10,000 examples), randomly coloring each digit either red or green, and adding these two colors as concepts. There are two tasks: The first is whether the digit is even, and the second is whether it is odd. For these tasks, the concepts related to color are irrelevant.
@@
This is a large-scale face attributes dataset with more than 200K celebrity images [liu2015faceattributes]. Each image has 40 concept annotations. As tasks, we take the concepts $Wavy\_Hair$, $Black\_Hair$, and $Male$, removing them from the concept set.
@@
In this dataset, the task is to predict bird species from images, where the concepts are 112 bird characteristics such as tail color, wing pattern, etc  [cub]. CUB originally consists of 200 tasks, of which we take the first 10. Some concepts are strongly imbalanced (some are True in only ±$\pm 0.5\%$ of examples, others in $40\%$, etc.) and there is a large task imbalance (each task is True in only ±$\pm 0.5\%$ of examples).
@@
This is a text-based dataset that consists of reviews, where the task is to classify them as positive or negative [abraham2022cebab]. There are 12 concepts (food, ambience, service, noise, each either unknown, bad or good) and 2 tasks (positive or negative review).
@@ C.2 Training
When departing from pure probabilistic semantics, CBMs allow not only the use of concepts as binary variables, but they allow for concepts to be passed to the task predictor together with their prediction scores, which is called employing soft concepts (vs hard concepts). While some CBMs use soft concepts as this results in higher task accuracy, the downside of this is that the use of soft concepts also comes with the introduction of input distribution leakage: The concept probabilities encode much more information than what is related to the concepts, severely harming the interpretability of the model [marconato2022glancenets, havasi2022addressing]. For this reason, in our experiments, all models use hard concepts, which is realized by thresholding the soft concept predictions at 50%.
@@
For MNIST+, MNIST+ and C-MNIST, we train directly on the images. For CelebA and CUB, instead of training on the images, we train the models on pretrained ResNet18 embeddings [he2016deep]. Specifically, using the torchvision library, we first resize the images to width and height 224 (using bi-linear interpolation), then normalize them per channel with means $(0.485,0.456,0.406)$ and standard-deviations $(0.229,0.224,0.225)$ (for CelebA only). Finally, we remove the last (classification) layer from the pretrained ResNet18 model, use the resulting model on each image, and flatten the output, resulting in an embedding. For CEBAB, we use a pretrained BERT model [devlin2018bert] to transform the input into embeddings. Specifically, using the transformers library [wolf-etal-2020-transformers], we create a BERT model for sequence classification from the pretrained model ’bert-base-uncased’ with 13 labels (1 per concept and 1 representing both tasks). Then, we finetune this model for 10 epochs with batch size 2, 500 warmup steps, weight decay 0.01 and 8 gradient accumulation steps. After training, we use this model to transform each example into an embedding by outputting the last hidden states for that example.